# Chinese Large Language Models via Unconstrained Generation thanks: * These authors contribute equally

### v2

@@ 5 Conclusion
Firstly, although we have utilized hallucination ranking, automatic labeling, human rechecking, and various other techniques mentioned in Appendix A to ensure the quality of data annotation, with over 5,000 data entries, there is still a possibility of labeling errors. We have mobilized the power of the open-source community to collectively improve our dataset. Additionally, the dataset creation process is flexible, allowing for dataset expansion into English and broader domains, such as mathematical reasoning and programming codes.
@@
Appendix A The UHGEval Dataset
A.1 Dive into Human Rechecking Process
@@
A.2 Analysis of Final Dataset
@@
A.3 An example from the UHGEval dataset
@@
Appendix B Experiments
B.1 LLMs Employed in This Research
@@ B.2 Evaluation Method
B.3 UHGEval Framework in Detail
@@
Appendix C More Related Works
C.1 Large Language Models
LLMs can be classified according to two dimensions. The first dimension concerns the openness of the model weights. For example, open-source models include Meta’s LLaMA Touvron et al. (2023), Tsinghua University’s GLM Du et al. (2022), and Alibaba’s Qwen Bai et al. (2023), while closed-source models feature OpenAI’s GPT OpenAI (2023), Baidu’s ERNIE Bot Sun et al. (2021), and Anthropic’s Claude [^6] https://www.anthropic.com/index/introducing-claude, among others. The second dimension differentiates between the use of a PLM or a supervised fine-tuned (SFT) model for specific inferences. A PLM is a language model trained on extensive unlabeled textual data to discern underlying patterns, structures, and semantic knowledge within the corpus. Conversely, an SFT model involves further training a PLM with labeled datasets tailored to a specific task, to improve performance in that area. Many open-source models, including LLaMA, GLM, and Qwen, have made their PLM weights publicly available. For SFT models, users can access the chat variants of open-source models or the API services provided by closed-source models. In our research, we focus primarily on evaluating closed-source GPT series models and open-source Chinese chat models.
C.2 Hallucinations in LLM
Appendix D Figures in Chinese
@@
Appendix E Prompt Templates
In these templates, the orange text represents intent and instruction, the green text represents demonstrations, and the black text represents specific questions.

### v3

@@ 5 Conclusion
Firstly, although we have utilized hallucination ranking, automatic labeling, human rechecking, and various other techniques mentioned in Appendix C to ensure the quality of data annotation, with over 5,000 data entries, there is still a possibility of labeling errors. We have mobilized the power of the open-source community to collectively improve our dataset. Secondly, the dataset creation process is flexible, allowing for dataset expansion into English and broader domains, such as mathematical reasoning and programming codes. Thirdly, a minor error in the dataset creation process has resulted in a relatively unbalanced distribution of the dataset across the five different LLMs used for generation. A detailed analysis of this issue can be found in Appendix G.
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 62072463, 71531012), the National Social Science Foundation of China (Grants No. 18ZDA309), the Research Seed Funds of the School of Interdisciplinary Studies at Renmin University of China, and the Opening Project of the State Key Laboratory of Digital Publishing Technology of the Founder Group.
Appendix A Comparisons with Other Datasets
Below are specific comparisons with other datasets and the significance of unconstrained generation.
A.1 TruthfulQA
The TruthfulQA dataset encompasses three modes of evaluation, with the primary mode being generative. In this mode, a problem is presented to the model, which then freely generates content that is assessed by humans or a fine-tuned GPT-judge. The other two modes are single- / multiple-choice questions. In these modes, a problem along with reference options is provided, the model makes a selection, and accuracy is calculated.
Figure 1 in the TruthfulQA paper includes statements indicating that some content is freely generated by GPT-3. This might be somewhat misleading. The content is used solely to evaluate the performance of the GPT-3 model in generative evaluation and is not part of the dataset. The actual free generation pertains to the "reference options" in the single- / multiple-choice questions. These reference options are manually crafted in TruthfulQA.
Appendix C of the paper details the method used to create the reference options:
Reference answers for each question in TruthfulQA are constructed as follows:
We take a set of true answers directly from Wikipedia (or the listed source). We then try to provide coverage of common variations on this answer…
We follow a similar process for generating false answers, but widen the answer set by running internet searches for [common misconceptions / superstitions / conspiracies around X] where relevant, as there tend to be many possible imitative false answers that are not always covered in a single source…
A.2 HaluEval
The problem types within this benchmark are all judgment questions, tasked with determining whether an option contains hallucinations. Accordingly, they also provide reference options. However, their method of generating these options is targeted. An example of how they generate options is: "You are trying to answer a question but misunderstand the question context and intention." They then take such generated texts and real texts, placing them together for downstream models to evaluate for the presence of hallucinations.
A.3 HaDes
HaDes evaluates a model’s ability to identify hallucinated words within a given text. However, the method used to generate these hallucinations involves randomly altering correct text, thereby transforming some words into hallucinations. This approach to generating errors leads to a distributional bias compared to the hallucinations that arise from the model’s free output.
In summary, most existing datasets related to hallucinations are purposefully and manually generated with constraints. They do not represent the hallucinations that might be collected while the model is addressing user queries or responding to users in real-world scenarios. This raises the question: Are the errors generated in this manner truly reflective of the mistakes a model would make? Hence, in creating our dataset, we allowed the model to output freely, collecting only those portions where hallucinations occurred. This represents one of the major challenges in our work.
A.4 Why Is Unconstrained Generation Important?
In datasets like TruthfulQA, HaluEval, and HaDes, it’s challenging to pinpoint exactly why a model might produce hallucinations. These texts, potentially containing inaccuracies, are designed to assess whether a downstream model can identify errors within a text. However, our dataset enables a genuine evaluation of model hallucinations, even tracing their origins. For instance, in our dataset, the entry with ID doc_000002 features hallucinations generated by the Baichuan2-13B model. The terms related to "economic development" and others, totaling five words, are involved in these hallucinations, while words like "China" and another set of five words are not. This distinction allows us to investigate whether there are differences in the token logits, the states of hidden layers, etc., between the words associated with hallucinations and those without, in the context of the Baichuan model. Theoretically analyzing the causes of hallucinations within the Baichuan model is part of our ongoing work. This approach is something that other benchmarks cannot offer, as their hallucinations are not freely produced by the model, and in some cases, not even generated by models.
Appendix B More Related Works
B.1 Large Language Models
LLMs can be classified according to two dimensions. The first dimension concerns the openness of the model weights. For example, open-source models include Meta’s LLaMA Touvron et al. (2023), Tsinghua University’s GLM Du et al. (2022), and Alibaba’s Qwen Bai et al. (2023), while closed-source models feature OpenAI’s GPT OpenAI (2023), Baidu’s ERNIE Bot Sun et al. (2021), and Anthropic’s Claude [^6] https://www.anthropic.com/index/introducing-claude, among others. The second dimension differentiates between the use of a PLM or a supervised fine-tuned (SFT) model for specific inferences Zhu et al. (2024). A PLM is a language model trained on extensive unlabeled textual data to discern underlying patterns, structures, and semantic knowledge within the corpus. Conversely, an SFT model involves further training a PLM with labeled datasets tailored to a specific task, to improve performance in that area. Many open-source models, including LLaMA, GLM, and Qwen, have made their PLM weights publicly available. For SFT models, users can access the chat variants of open-source models or the API services provided by closed-source models. In our research, we focus primarily on evaluating closed-source GPT series models and open-source Chinese SFT models.
B.2 Hallucinations in LLM
Appendix C The UHGEval Dataset
C.1 Dive into Human Rechecking Process
@@
Annotation Team
Our annotators are all Chinese nationals with Chinese as their native language, each holding at least a Master’s degree in Journalism. We collaborated with a well-known, sizable news organization in China, Xinhua News Agency. Some of their staff joined our research team and participated in data annotation for this project. There were a total of 9 annotators involved in this project, with a gender ratio of 1:2 (male to female). Regarding their compensation, they first received a standard employee salary. Additionally, they were paid an extra 3 RMB for each data item annotated, with each item taking about 40 seconds to annotate. Besides the annotators, our engineering team and experts from the journalism industry at Xinhua News Agency participated in the data review process, totaling 3 people. Our main responsibility was to supervise and review the quality of the annotations. The entire annotation process lasted for 22 days.
C.2 Analysis of the Final Dataset
@@
C.3 Data Volume for Each Step
In this section, we present the data volume at various stages of our dataset creation process for reference and transparency.
  • Data volume of original news dataset: 737,766
  • Data volume after preprocessing: 25,005 (filtering out outliers in dimensions such as length and news type)
  • Data volume after generating candidate hallucination text: 17,503 (filtering out data items that did not generate appropriate continuations, such as those with excessively short length or insufficient extracted keywords)
  • Data volume with hallucination in machine-labeled text: 8,314 (filtering out texts deemed by the machine to lack hallucination)
  • Data volume after human annotator labeling: 5,141 (filtering out instances not verified as hallucination upon manual review, or deemed inappropriate, such as those repeating previous content or generating text no longer of news type but rather comprehension questions)
C.4 An example from the UHGEval Dataset
@@
Appendix D Experiments
D.1 LLMs Employed in This Research
@@ D.2 Evaluation Method
D.3 UHGEval Framework in Detail
@@
Appendix E Figures in Chinese
@@
Appendix F Prompt Templates
In these templates, the orange text represents intent and instruction, the green text represents demonstrations, and the black text represents specific questions. The template may be very long, and we may use ellipses to omit some content in the middle. The original templates are in Chinese, and we also provide English translations.
Figure 13: Prompt template for extracting keywords
@@ Appendix F Prompt Templates
Appendix G Clarification of Imbalance of the Dataset in Term of Models
G.1 Imbalance Phenomenon
The distribution of the final hallucination dataset in terms of five different LLMs used in the generation is shown in Table 7.
Generated by Count Share
Baichuan 3425 66.62%
ChatGLM 327 6.36%
Xinyu 424 8.25%
InternLM 487 9.47%
Qwen 478 9.30%
Table 7: Dataset distribution by LLMs
You may notice a disproportionately high representation of the Baichuan model, a discrepancy linked to an oversight at the outset. Initially, the generation of candidates did not fully leverage the available models, primarily because the Baichuan models exhibited superior instruction-following capabilities for data generation. Consequently, we solely utilized five instances of the Baichuan model for ranking to generate candidate data items. It was only later that four additional models were incorporated, employing a total of five distinct model instances for ranking to generate candidate data items. This approach resulted in a final dataset that lacks a relatively balanced distribution among the different models.
G.2 Does the Imbalance Lead to Unreliable Outcomes?
A quick answer is: No. Here are the justifications.
We conducted a simple empirical study. Since the generated hallucinated texts are only used in selective and discriminative evaluations, if the imbalance significantly affects the experimental outcomes, it would only impact these two types of assessments. Fortunately, we have saved every piece of intermediate experimental results, allowing us to uniformly sample those results across models and re-aggregate the results to observe changes in key metrics like average accuracy. Specifically, we uniformly and randomly sampled 327 data points for each model and used a combined dataset of 327*5=1635 data points to re-aggregate the results. Below is a comparison of the experimental outcomes. Table 8 and Table 9 present the original results and the new results, respectively, while Table 10 displays the differences between them.
In analyzing the differences, we calculated three metrics to illustrate the magnitude of change. The average change in the three columns of accuracy metrics is approximately -0.0147, with a standard deviation of 0.0150. Furthermore, the average absolute change in their rankings is 0.606 (because only six pairs of closely ranked models undergo internal swaps). Lastly, upon separately examining the Baichuan model, which was anticipated to be most affected, we found its change to be not the greatest. Therefore, we can assert that the variance in model proportions has an insignificant impact on the outcomes.
Moreover, intuitively, since all these models are Chinese text generation models and the content they generate was also reviewed by our manual annotators, we eliminated some data items that significantly deviated the text type from the mean during annotation (for example, repetitions of previous content, or continuations that diverged from news to generating a reading comprehension question, etc.). This enhanced the overall consistency of the final dataset, making the task of distinguishing hallucinations in the experiment unrelated to the source of those hallucinations.
Discriminative-Keyword Discriminative-Sentence Selective
avg.acc. avg.#kws #valid avg.acc. #valid acc. #valid
Aquila-34B 53.62% 3.00 3719 49.86% 5009 54.29% 4319
Baichuan2-13B 51.63% 3.13 4478 46.88% 5047 50.23% 5130
Baichuan2-53B 52.13% 2.98 1656 50.81% 1478 54.67% 4443
ChatGLM2-6B 50.80% 3.10 4289 43.87% 5130 43.59% 5130
GPT3.5-Turbo 53.72% 3.08 4183 50.02% 5039 49.03% 5103
GPT4-0613 70.04% 3.07 4100 57.42% 5024 55.20% 5047
GPT4-1106 69.48% 3.10 4189 57.38% 4903 60.35% 4752
InternLM-20B 50.92% 3.10 4388 51.01% 5130 49.43% 5130
Qwen-14B 52.86% 3.13 4478 50.58% 5130 54.74% 5130
Xinyu-7B 49.58% 3.12 4451 48.66% 5014 50.58% 5130
Xinyu2-70B 52.94% 3.12 4482 55.04% 5128 57.93% 5129
Table 8: Original Results
Discriminative-Keyword Discriminative-Sentence Selective
avg.acc. avg.#kws #valid avg.acc. #valid acc. #valid
Aquila-34B 54.17% 3.18 1178 50.98% 1582 57.34% 1362
Baichuan2-13B 51.61% 3.29 1398 50.34% 1608 51.93% 1629
Baichuan2-53B 52.68% 3.06 525 51.46% 479 56.70% 1432
ChatGLM2-6B 51.27% 3.27 1357 47.02% 1629 46.04% 1629
GPT3.5-Turbo 54.38% 3.23 1291 51.87% 1601 50.06% 1620
GPT4-0613 70.03% 3.23 1277 59.73% 1593 58.79% 1582
GPT4-1106 68.24% 3.25 1305 61.28% 1547 65.34% 1503
InternLM-20B 51.06% 3.23 1348 52.42% 1629 53.53% 1629
Qwen-14B 53.84% 3.29 1404 51.20% 1629 53.96% 1629
Xinyu-7B 49.51% 3.29 1389 48.74% 1582 50.58% 1629
Xinyu2-70B 54.30% 3.29 1402 58.24% 1627 59.28% 1628
Table 9: New Results
Discriminative-Keyword Discriminative-Sentence Selective
avg. acc. avg. #kws #valid avg. acc. #valid acc. #valid
Aquila-34B -0.55% -0.18 2541 -1.12% 3427 -3.05% 2957
Baichuan2-13B 0.02% -0.16 3080 -3.46% 3439 -1.70% 3501
Baichuan2-53B -0.55% -0.08 1131 -0.65% 999 -2.03% 3011
ChatGLM2-6B -0.48% -0.17 2932 -3.15% 3501 -2.45% 3501
GPT3.5-Turbo -0.67% -0.15 2892 -1.85% 3438 -1.03% 3483
GPT4-0613 0.01% -0.16 2823 -2.31% 3431 -3.59% 3465
GPT4-1106 1.24% -0.14 2884 -3.90% 3356 -4.98% 3249
InternLM-20B -0.14% -0.13 3040 -1.41% 3501 -4.10% 3501
Qwen-14B -0.98% -0.17 3074 -0.62% 3501 0.78% 3501
Xinyu-7B 0.07% -0.17 3062 -0.07% 3432 0.00% 3501
Xinyu2-70B -1.36% -0.16 3080 -3.20% 3501 -1.35% 3501
Table 10: Difference (Original - New)